6. A solid iron pole consists of a cylinder of height 220 cm and base diameter 24 cm, which is surmounted by another cylinder of height 60 cm and radius 8 cm. Find the mass of the pole, given that 1 cm3 of iron has approximately 8 g mass.

Grade:12th pass

1 Answers

Pawan Prajapati
2 years ago
Given, the height of the big cylinder (H) = 220 cm Radius of the base (R) = 24/12 = 12 cm So, the volume of the big cylinder = πR2H = π(12)2 × 220 cm3 = 99565.8 cm3 Now, the height of smaller cylinder (h) = 60 cm Radius of the base (r) = 8 cm So, the volume of the smaller cylinder = πr2h = π(8)2×60 cm3 = 12068.5 cm3 ∴ Volume of iron = Volume of the big cylinder+ Volume of the small cylinder = 99565.8 + 12068.5 =111634.5 cm3 We know, Mass = Density x volume So, mass of the pole = 8×111634.5 = 893 Kg (approx.)